OpenStudy (anonymous):

HELP!!!!!(Question) The expression 4n-1 is a rule to describe a certain sequence. What are the first tree terms of the sequence. A 3, 7, 11 B 4, 8 , 12 C 3, 6, 9 D 4, 7, 11

OpenStudy (cwrw238):

n = 1 we get 4(1) - 1 = 3 n = 2 ... .... 4(2) - 1 = 7 can you say which one it is?

OpenStudy (anonymous):

Hmm B? maybe?

OpenStudy (cwrw238):

no - i've given you the first 2 terms so ..

OpenStudy (cwrw238):

the formula for each term is 4(n) - 1 so first term when n = 1 is obtained by replacing n by 1 - so its 4(1) - 1 = 3

OpenStudy (cwrw238):

yw